Circular No. 8529

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V378 SERPENTIS AND V2361 CYGNI D. Gaggero, I. Martire, R. Poggiani, V. Puccetti, S. N. Shore, and

E. Tognelli, Department of Physics, University of Pisa; and S. Bernabei, Osservatorio Astrofisica, Bologna, and Istituto Nazionale di Astrofisica (INAF), report 0.4–1.0-μm observations of V378 Ser and V2361 Cyg using the Loiano 1.52-m telescope (+ BFOSC grism spectrograph) of the Bologna Observatory/INAF during Apr. 29.8–May 2.2 UT (at resolutions of 700 and 900). V378 Ser (IAUC 8505) showed weak Hα P-Cyg absorption (Hβ and Hγ indicate a Balmer progression in emission intensity, while the velocity of the minimum of each of the three Balmer absorption troughs was −2200 km/s); stronger absorption is seen for Paschen 10–13, and weak emission appears at C I 908.5- and 940.6-nm and O I 777.3- and 844.2-nm but with similar velocities. No He II or O II is observed. Strong P-Cyg absorption is detected for He I 587.5-nm at −1500 km/s and also for Fe II (42, 48, 49). The spectrum resembles those of the nova in the Large Magellanic Cloud 1988 No. 1 and V443 Sct at two weeks after maximum — consistent with the ejecta being in the pre-nebular phase (no change observed in line profile or intensity during the 3 days of observation). The spectrum is developing relatively slowly for the observed optical velocities. V2361 Cyg (IAUC 8483) displays strong emission at Hα with principal asymmetric peaks at −700 km/s (weaker) and +705 km/s, and a still-weaker third peak at −200 km/s (FWHM = 2200 km/s and FWZI = 4500 km/s for the entire profile); O I 844.2-nm shows the same peak velocities and intensities as that of the Hα emission (FWHM = 2000 km/s and FWZI 3800 km/s); the Paschen series has similar widths; and other emission lines (e.g., [O III] 436.3-, 495.9-, and 500.7-nm; He I 587.5-nm; O I 777.3, 844.2 nm; N I 821.6-nm; and Mg II 823.2-nm) show symmetric peaks at the same peak velocities. The Hα/Hβ intensity ratio is consistent with the high reported reddening (IAUC 8524). Both novae resemble CO-type outbursts.

COMETS C/2005 F3 AND C/2005 F4 (SOHO) Additional diffuse Kreutz sungrazers (cf. IAUC 8527); C/2005 F3 had

no tail (too faint for photometry, but at mag ∼ 8), while C/2005 F4 had a possible hint of a tail and reached mag ∼ 6.5 at 7.1R_{⊙} (Mar. 28.842 UT).

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 F3 Mar. 28.017 0^{h}35.^{m}5 + 2^{o}17′ C2  MM 2005-J25
C/2005 F4  28.429 0 44.9 + 2 08 C3/2 HS 2005-J25
2005 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green